# Bayesian power spectral density estimation for LISA noise based on P-splines with a parametric boost

### v1

@@ Abstract
Flexible and efficient noise characterization is crucial for the precise estimation of gravitational wave parameters. We introduce a fast and accurate Bayesian method for estimating the power spectral density (PSD) of long, stationary time series tailored specifically for LISA data analysis. Our approach models the PSD as a geometric mean of a parametric and a nonparametric component, combining the computational efficiency of parametric models with the flexibility to capture deviations from theoretical expectations. The nonparametric component is expressed by a mixture of penalized B-splines. Adaptive, data-driven knot placement performed once during initialization eliminates computationally expensive reversible-jump Markov Chain Monte Carlo, while hierarchical roughness penalty priors prevent overfitting. This design yields stable, flexible PSD estimates with runtimes of minutes instead of hours. Validation on simulated autoregressive AR(4) data demonstrates estimator consistency. It shows that well-matched parametric components reduce the integrated absolute error compared to an uninformative baseline, requiring fewer spline knots to achieve comparable accuracy. Applied to a year of simulated LISA $X$-channel noise, our method achieves relative integrated absolute errors of $\mathcal{O}(10^{-2})$ with computation times less than three minutes, which makes it suitable for iterative analysis pipelines and multi-year mission datasets.
@@ I Introduction
The space-based Laser Interferometer Space Antenna (LISA) mission will open the low-frequency mHz band of gravitational wave astronomy, observing sources ranging from Galactic white-dwarf binaries to mergers of supermassive black holes [1]. Unlike ground-based detectors, LISA will observe continuously for years (aside from occasional data gaps), with signals from multiple sources overlapping throughout the data stream. This continuous observation eliminates the possibility of “quiet” periods for off-source noise estimation [2], and LISA’s unique noise characteristics—including instrumental noise, acceleration noise, and a dominant Galactic foreground—combined with multi-year observation times, demand computationally efficient methods capable of handling extremely long time series. For context, a full LISA mission (4-6 years) sampled at 1 Hz would generate $\sim 10^{8}$ observations, making computational efficiency paramount. In ground-based detectors such as LIGO–Virgo–KAGRA, noise is typically estimated from off-source segments using the Welch method [3], averaging periodograms computed from multiple signal-free segments to obtain a consistent power spectral density (PSD) estimate [4]. This PSD can be used as a plug-in estimate in the Whittle likelihood, enabling fast, low-latency analyses. However, for LISA and other next-generation detectors such as the Einstein Telescope (ET) [5] and Cosmic Explorer (CE) [6], the absence of off-source, noise-only segments renders these approaches infeasible, motivating flexible, on-source PSD estimation methods that are both accurate and computationally efficient.
@@
Accurate modeling of the detector PSD is critical for unbiased parameter estimation and well-calibrated uncertainties, since misestimating the noise spectrum can lead to biased parameter estimates and over-confident credible intervals [7]. Several algorithms have been developed for flexible PSD estimation, particularly for LISA. BayesWave [8] uses a reversible-jump Markov Chain Monte Carlo (RJMCMC) [9, 10] algorithm to fit a trans-dimensional model with a Morlet-Gabor wavelet frame, simultaneously estimating the gravitational wave signal and the noise PSD via BayesLine [11] using Lorentzians and splines. Frequency-binned approaches such as SGWBinner [12] fit power laws in discrete bins, while Gaussian process models for the log-PSD [13] provide weakly parametric smoothing. Several works model deviations from fixed parametric PSDs using B-splines or AKIMA splines with unknown numbers and locations of knots [14, 15, 16], again relying on RJMCMC, and some incorporate parametric corrections to improve estimation of sharp spectral features. Alternative Bayesian nonparametric methods [17, 18] avoid RJMCMC by inducing spline coefficients through a Dirichlet process prior on a cumulative distribution function. In particular, Kirch et al. (2019) [17] further improve the estimation of sharp PSD peaks by employing a nonparametrically corrected likelihood that combines autoregressive time-domain modeling with frequency-domain corrections. While flexible, Markov Chain Monte Carlo (MCMC) based methods remain computationally expensive, particularly due to repeated transformations between time and frequency domains, highlighting the need for scalable alternatives for multi-year datasets.
@@
Penalized splines (P-splines) provide a computationally efficient solution [19, 20]. By using a deliberately large B-spline basis with a hierarchical roughness penalty prior, P-splines allow flexible smoothing while controlling overfitting.
@@
This work presents a novel log-P-spline algorithm tailored to LISA’s computational demands. Defining the basis on a logarithmic frequency grid offers distinct advantages: the basis is first evaluated on a uniform grid in normalized space $[0,1]$ and then mapped logarithmically to the physical frequency range $f_{\min},f_{\rm max}]$. This mapping produces denser coverage at low frequencies, where LISA sensitivity is highest and spectral features such as $1/f$ acceleration noise, Galactic confusion noise, and instrumental artefacts are most prominent, while avoiding overparameterisation at high frequencies. Logarithmic spacing is consistent with the standard log–log representation of spectral densities and ensures that the roughness penalty acts approximately uniformly across frequency decades. This logarithmic parameterization is thus ideally suited to LISA’s noise characteristics and scientific requirements. Modeling the log-PSD removes the positivity constraint on the spline coefficients and allows the roughness penalty to operate uniformly across frequency decades. Our approach uses a corrected likelihood exclusively in the frequency domain, avoiding the costly time-frequency transformations required by methods such as Kirch et al. (2019) [17]. The PSD is modeled as a geometric mean of parametric and nonparametric components, allowing for flexible correction of potentially misspecified parametric models while retaining the speed and flexibility of P-splines. The parametric component, incorporating known LISA characteristics such as acceleration noise scaling and instrumental transfer functions, further enhances the estimation of sharp spectral features. The algorithm employs a blocked Whittle likelihood to handle extensive time series, dividing the dataset into segments with a shared PSD. We demonstrate the method on a one-year-long simulated LISA time series sampled at 1 Hz ($\sim 3.15\times 10^{7}$ observations), with computation times less than three minutes.
@@ II P-spline Model for the Log PSD
where $\Gamma_{j}(h)=\left(\gamma_{lm}(h)\right)=\mathbb{E}(Z^{j}_{t}Z^{j}_{t+h})$, with $h$ being the lag, is the Fourier transform of the auto-covariance function. This leads to the widely used Whittle likelihood function [21], particularly prevalent in gravitational wave research [22]. Under the additional assumptions of segment independence and identical PSDs $S(f_{l})=S_{j}(f_{l})$ for $j=1,\ldots,J$, the blocked Whittle likelihood is obtained as
@@
$\propto\exp\left\{-J\sum_{l=1}^{v}\left(\log(S(f_{l}))-\frac{\bar{I}(f_{l})}{S(f_{l})}\right)\right\},$ (3)
@@
To efficiently account for expected spectral features, we include a parametric PSD, $S_{\text{par}}(f)$, representing these components of the spectrum within the likelihood. The full spectral density is then modeled as the geometric mean of this parametric component and a nonparametric P-spline component, $S_{\text{npar}}(f)$, allowing flexibility to capture any deviations while retaining the overall structure suggested by $S_{\text{par}}(f)$, and is expressed as
@@
Equations (57) show that the geometric mean model turns out to be equivalent to modeling a correction of a parametric model as in Equation (13) of [14] where the parametric PSD is taken to be that of some instrumental design specification.
@@
where B-spline densities $b_{k,r}$ are normalized B-spline functions, i.e., integrate to $1$, of fixed degree $r$ with the knot sequence $\xi$ which is fixed on the interval of Fourier frequencies at which our periodogram is defined, and $\boldsymbol{\lambda}$ is the parameter vector comprising $\lambda_{1},\lambda_{2},\dots,\lambda_{K}\in\mathbb{R}$.
@@
Our method employs a B-spline basis defined on a logarithmically spaced frequency grid, which offers several advantages over the more conventional linearly spaced basis in PSD estimation. This non-uniform spacing allows the model to capture such features with higher fidelity, while avoiding overparameterization of the high-frequency regime, where, for example, the LISA instrument is less sensitive. Furthermore, logarithmic spacing is consistent with the standard visual and statistical treatment of spectral densities, which are frequently examined on a log–log scale.
@@
The direct modeling of the logarithm of the PSD results in unconstrained values of $\lambda$, which stands in contrast to the approach in [19], where the weights were explicitly constrained to sum to one, thereby introducing an additional layer of complexity. The proposed formulation simplifies the MCMC sampler by removing this restriction and enhancing its computational efficiency.
@@
B-spline functions require a predetermined set of knots. When their number is allowed to vary, methods such as RJMCMC or the stick-breaking representation of the Dirichlet process [18] become necessary. Both approaches are challenging to tune, complicate the assessment of Markov chain convergence, and substantially increase computational cost. We select a large but fixed number of B-splines with predetermined knots in our proposal. Although this choice may naturally increase the risk of overfitting, it is mitigated by the penalty on the B-spline coefficients imposed through their prior distribution.
@@
As shown in [19], judicious selection of the knot locations allows P-splines to retain the flexibility of B-splines while improving computational efficiency. We implement two options for placing the knots of the B-spline basis functions in our models. In the first approach, we place equidistant knots on the logarithmic scale, analogous to the traditional P-spline methods that employ equally spaced knots. In the second approach, the quantile-based knot scheme proposed in [19] is applied to the ratio of the averaged periodogram and the parametric PSD ($\bar{I}/S_{\text{par}}(f_{l})$). In this way, more knots are allocated in regions where the parametric model fails to capture the pattern of the periodogram.
@@
If the $\lambda$ coefficients were allowed to vary too freely, the resulting representation of the log correction would fluctuate excessively and lead to overfitting. Therefore, in the frequentist approach, one adds a penalty term to the likelihood function, e.g., $\phi(\Delta\lambda_{k})=\phi(\lambda_{k}-\lambda_{k-1})$ or $\phi(\Delta^{2}\lambda_{k})$ that penalizes the first or second order differences $\Delta^{2}\lambda_{k}=(\lambda_{k}-\lambda_{k-1})-(\lambda_{k-1}-\lambda_{k-2})$, respectively. Within the Bayesian context, instead of minimizing the penalized log-likelihood, the penalty term is naturally included in the prior distribution of the coefficients.
To avoid dependence on the choice of the penalty parameter $\phi$, a hierarchical prior structure is used as follows
$\boldsymbol{\lambda}|\phi\sim\mathcal{N}_{K}(\textbf{0},(\phi\textbf{P})^{-1})$ (9)
$\phi|\delta\sim\mathrm{Gamma}(\alpha_{\phi},\delta\beta_{\phi})$ (10)
$\delta\sim\mathrm{Gamma}(\alpha_{\delta},\beta_{\delta})$ (11)
where $\textbf{P}=\textbf{D}^{\top}\textbf{D}+\epsilon\textbf{I}_{K}$ is the penalty matrix–full rank matrix for any small quantity $\epsilon$, e.g., $10^{-6}$–with D denoting the difference matrix. For non-equidistant knots, we construct D using the derivative-based penalties described by [23] and used in [19]. In our applications, we use first-order penalties and third-order B-spline basis functions.The rate parameters $\beta_{\phi}$ and $\beta_{\delta}$ are given small prior values, e.g., $10^{-4}$, and the shape parameters $\alpha_{\phi}$ and $\alpha_{\delta}$ values are around 1 for robustness.
@@
We use a blocked Gibbs sampler to sample from the joint posterior $p(\boldsymbol{\theta}|\textbf{Y})$ by cyclically sampling from each of the full conditional posterior distributions. $\phi$ and $\delta$ can be sampled directly from their full conditional posterior distributions
@@
Within the Gibbs-sampling framework, $\boldsymbol{\lambda}$ is sampled either using a Metropolis–Hastings method for component-wise updates [19] or an adaptive Metropolis–Hastings (AMH) algorithm for joint updates [24].
For computational efficiency, the knots are placed at the beginning of the algorithm and remain fixed throughout. To ensure sufficient flexibility, we can use a large number of knots, as illustrated in the Application section, so that the model can capture the underlying variations of the PSD. Additionally, the penalty prior helps guard against overfitting.

### v2

@@ Abstract
Flexible and accurate noise characterization is crucial for the precise estimation of gravitational-wave parameters. We introduce a Bayesian method for estimating the power spectral density (PSD) of long, stationary time series, explicitly tailored for LISA data analysis. Our approach models the PSD as the geometric mean of a parametric and a nonparametric component, combining the knowledge from parametric models with the flexibility to capture deviations from theoretical expectations. The nonparametric component is expressed by a mixture of penalized B-splines. Adaptive, data-driven knot placement, performed once at initialization, removes the need for reversible-jump Markov chain Monte Carlo, while hierarchical roughness-penalty priors prevent overfitting. Validation on simulated autoregressive AR(4) data demonstrates estimator consistency and shows that well-matched parametric components reduce the integrated absolute error compared to an uninformative baseline, requiring fewer spline knots to achieve comparable accuracy. Applied to one year of simulated LISA X-channel (univariate) noise, our method achieves relative integrated absolute errors of $\mathcal{O}(10^{-2})$, making it suitable for iterative analysis pipelines and multi-year mission data sets.
@@ I Introduction
The space-based Laser Interferometer Space Antenna (LISA) mission will open the low-frequency mHz band of gravitational wave astronomy, observing sources ranging from Galactic white-dwarf binaries to mergers of supermassive black holes [amaro-seoane_laser_2017]. Unlike ground-based detectors, LISA will observe continuously for years (aside from occasional data gaps), with signals from multiple sources overlapping throughout the data stream. This continuous observation eliminates the possibility of “quiet” periods for off-source noise estimation [littenberg_prototype_2023], and LISA’s unique noise characteristics—including instrumental noise, acceleration noise, and a dominant Galactic foreground—combined with multi-year observation times, demand computationally efficient methods capable of handling extremely long time series. For context, a full LISA mission (4-6 years) sampled at 1 Hz would generate $\sim 10^{8}$ observations, making computational efficiency paramount. In ground-based detectors such as LIGO–Virgo–KAGRA, noise is typically estimated from off-source segments using the Welch method [welch_use_1967], averaging periodograms computed from multiple signal-free segments to obtain a consistent power spectral density (PSD) estimate [abbott_guide_2020]. This PSD can be used as a plug-in estimate in the Whittle likelihood, enabling fast, low-latency analyses. However, for LISA and other next-generation detectors such as the Einstein Telescope (ET) [punturo_einstein_2010] and Cosmic Explorer (CE) [reitze_cosmic_2019], the absence of off-source, noise-only segments renders these approaches infeasible, motivating flexible, on-source PSD estimation methods that are both accurate and computationally efficient.
@@
Accurate modeling of the detector PSD is critical for unbiased parameter estimation and well-calibrated uncertainties, since misestimating the noise spectrum can lead to biased parameter estimates and over-confident credible intervals [kirch_asymptotic_2024]. Several algorithms have been developed for flexible PSD estimation, particularly for LISA. BayesWave [cornish_bayeswave_2015] uses a reversible-jump Markov Chain Monte Carlo (RJMCMC) [Green:1995mxx, Umstatter:2005jd] algorithm to fit a trans-dimensional model with a Morlet-Gabor wavelet frame, simultaneously estimating the gravitational wave signal and the noise PSD via BayesLine [biscoveanu_quantifying_2020] using Lorentzians and splines. Frequency-binned approaches such as SGWBinner [flauger_improved_2021] fit power laws in discrete bins, while Gaussian process models for the log-PSD [PhysRevD.109.083029] provide weakly parametric smoothing. Several works model deviations from fixed parametric PSDs using B-splines or AKIMA splines with unknown numbers and locations of knots [muratore_impact_2023, santini_flexible_2025, baghi_uncovering_2023], again relying on RJMCMC, and some incorporate parametric corrections to improve estimation of sharp spectral features. Alternative Bayesian nonparametric methods [kirch_beyond_2019, edwards_bayesian_2019] avoid RJMCMC by inducing spline coefficients through a Dirichlet process prior on a cumulative distribution function. In particular, Kirch et al. (2019) [kirch_beyond_2019] further improve the estimation of sharp PSD peaks by employing a nonparametrically corrected likelihood that combines autoregressive time-domain modeling with frequency-domain corrections. While flexible, Markov Chain Monte Carlo (MCMC) based methods remain computationally expensive, particularly due to repeated transformations between time and frequency domains, highlighting the need for scalable alternatives for multi-year datasets.
@@
Penalized splines (P-splines) provide a computationally efficient solution [maturana-russel_bayesian_2021, Eilers_Marx_Durbán_2015]. By using a deliberately large B-spline basis with a hierarchical roughness penalty prior, P-splines allow flexible smoothing while controlling overfitting.
@@
This work presents a novel log-P-spline algorithm tailored to LISA’s computational demands. For LISA applications, we use knots that are uniformly spaced on a logarithmic scale. This produces denser coverage at low frequencies, where LISA sensitivity is highest and spectral features such as $1/f$ acceleration noise, Galactic confusion noise, and instrumental artefacts are most prominent, while avoiding overparameterisation at high frequencies. Logarithmic spacing is consistent with the standard log–log representation of spectral densities and ensures that the roughness penalty acts approximately uniformly across frequency decades. This logarithmic parameterization is thus ideally suited to LISA’s noise characteristics and scientific requirements. Modeling the log-PSD removes the positivity constraint on the spline coefficients and allows the roughness penalty to operate uniformly across frequency decades. Our approach uses a corrected likelihood exclusively in the frequency domain, avoiding the costly time-frequency transformations required by methods such as Kirch et al. (2019) [kirch_beyond_2019]. The PSD is modeled as a geometric mean of parametric and nonparametric components, allowing for flexible correction of potentially misspecified parametric models while retaining the speed and flexibility of P-splines. The parametric component, incorporating known LISA characteristics such as acceleration noise scaling and instrumental transfer functions, further enhances the estimation of sharp spectral features. The algorithm employs a blocked Whittle likelihood to handle extensive time series, dividing the dataset into segments with a shared PSD. We demonstrate the method on a one-year-long simulated LISA time series sampled at 1 Hz ($\sim 3.15\times 10^{7}$ observations), with computation times less than three minutes.
@@ II P-spline Model for the Log PSD
where $\Gamma_{j}(h)=\left(\gamma_{lm}(h)\right)=\mathbb{E}(Z^{j}_{t}Z^{j}_{t+h})$, with $h$ being the lag, is the auto-covariance function. This leads to the widely used Whittle likelihood function [whittle_curve_1957], particularly prevalent in gravitational wave research [christensen_parameter_2022]. Under the additional assumptions of segment independence and identical PSDs $S(f_{l})=S_{j}(f_{l})$ for $j=1,\ldots,J$, the blocked Whittle likelihood is obtained as
@@
$\propto\exp\left\{-J\sum_{l=1}^{v}\left(\log(S(f_{l}))+\frac{\bar{I}(f_{l})}{S(f_{l})}\right)\right\},$ (3)
@@
To efficiently account for expected spectral features, we include a parametric PSD, $S_{\text{par}}(f,\boldsymbol{\theta}_{\text{par}})$, representing these components of the spectrum within the likelihood. The full spectral density is then modeled as the geometric mean of this parametric component and a nonparametric P-spline component, $S_{\text{npar}}(f,\boldsymbol{\theta}_{\text{npar}})$, allowing flexibility to capture any deviations while retaining the overall structure suggested by $S_{\text{par}}(f,\boldsymbol{\theta}_{\text{par}})$. For simplicity, we refer to $S_{\text{par}}(f,\boldsymbol{\theta}_{\text{par}})$ and $S_{\text{npar}}(f,\boldsymbol{\theta}_{\text{npar}})$ as $S_{\text{par}}(f)$ and $S_{\text{npar}}(f)$, respectively. The full spectrum is therefore expressed as
@@
Equations (57) show that the geometric mean model turns out to be equivalent to modeling a correction of a parametric model as in Equation (13) of [muratore_impact_2023] where the parametric PSD is taken to be that of some instrumental design specification.
@@
where B-spline densities $b_{k,r}$ are normalized B-spline functions, i.e., integrate to $1$, of fixed degree $r$. $\boldsymbol{\xi}$ is the knot sequence $\boldsymbol{\xi}=\{f_{min}=\xi_{0}=\xi_{1}=\xi_{r},\leq\xi_{r+1}\leq\dots\leq\xi_{K}=\xi_{K+1}=\dots=\xi_{K+r}=f_{max}\}$, which is fixed on the interval of frequencies ($f_{min},f_{max}$) at which our periodogram is defined. Here, $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{K})^{\top}$ denotes the vector of spline coefficients, and $K$ is the total number of basis elements. The number of internal knots, which in our definition includes the boundary knots and is denoted $N_{\xi}$, is related to the number of basis functions by $K=N_{\xi}+r-1$, where $r$ is the spline degree (e.g., $r=3$ for cubic B-splines).
@@
Our method employs a B-spline basis defined on logarithmically spaced knots for noise spectral density estimation of LISA. Logarithmic spacing is consistent with the standard visual and statistical treatment of spectral densities, which are frequently examined on a log–log scale.
@@
The direct modeling of the logarithm of the PSD results in unconstrained values of $\lambda$, which stands in contrast to the approach in [maturana-russel_bayesian_2021], where the weights were explicitly constrained to sum to one, thereby introducing an additional layer of complexity. The proposed formulation simplifies the MCMC sampler by removing this restriction and enhancing its computational efficiency.
@@
B-spline functions require a predetermined set of knots. When their number is allowed to vary, methods such as RJMCMC or the stick-breaking representation of the Dirichlet process [edwards_bayesian_2019] become necessary. Both approaches are challenging to tune, complicate the assessment of Markov chain convergence, and substantially increase computational cost. We select a large but fixed number of B-splines with predetermined knots in our proposal. Although this choice may naturally increase the risk of overfitting, it is mitigated by the penalty on the B-spline coefficients imposed through their prior distribution.
@@
As shown in [maturana-russel_bayesian_2021], judicious selection of the knot locations allows P-splines to retain the flexibility of B-splines while improving computational efficiency. We implement two options for placing the knots of the B-spline basis functions in our models. In the first approach, we place equidistant knots on the logarithmic scale, analogous to the traditional P-spline methods that employ equally spaced knots. In the second approach, the quantile-based knot scheme from [maturana-russel_bayesian_2021] is applied to the ratio of the averaged periodogram and the parametric PSD ($\bar{I}/S_{\text{par}}(f_{l})$). This scheme allocates spline knots according to the empirical quantiles of this ratio, concentrating more knots in frequency regions where the periodogram exhibits strong deviations from the parametric model. Practically, the ratio is square-root transformed, standardized, and treated as a probability mass function whose cumulative distribution function is interpolated to define quantile locations. The resulting knot vector is fixed at initialization, so the computational cost of this step is negligible. This adaptive allocation enhances flexibility in regions where the parametric component underfits, while retaining smoothness elsewhere (for a more detailed discussion we refer the reader to [maturana-russel_bayesian_2021] and [Ruppert2003]). For clarity, throughout this work “smoothness of the spectra” refers to the smoothness of the ratio between the periodogram and the parametric model, since the P-spline prior is applied to the quantity $\log(c(f))$.
@@
If the $\lambda$ coefficients were allowed to vary too freely, the resulting representation of the log correction would fluctuate excessively and lead to overfitting. Therefore, in the frequentist approach, one adds a penalty term to the likelihood function, e.g., $\phi(\Delta\lambda_{k})=\phi(\lambda_{k}-\lambda_{k-1})$ or $\phi(\Delta^{2}\lambda_{k})$ that penalizes the first or second order differences $\Delta^{2}\lambda_{k}=(\lambda_{k}-\lambda_{k-1})-(\lambda_{k-1}-\lambda_{k-2})$, respectively. Within the Bayesian context, instead of minimizing the penalized log-likelihood, the penalty term is naturally included in the prior distribution of the coefficients. To avoid dependence on the choice of the penalty parameter $\phi$, a hierarchical prior structure is used as follows:
$\boldsymbol{\lambda}\mid\phi\sim\mathcal{N}_{K}(\mathbf{0},(\phi\mathbf{P})^{-1}),$ (9)
$\phi\mid\delta\sim\mathrm{Gamma}(\alpha_{\phi},\delta\beta_{\phi}),$ (10)
$\delta\sim\mathrm{Gamma}(\alpha_{\delta},\beta_{\delta}),$ (11)
where $\mathbf{P}$ is a full-rank penalty matrix. For non-equidistant knots, $\mathbf{P}$ is constructed using the derivative-based penalisation described by [wood_p-splines_2017] and used in [maturana-russel_bayesian_2021]. More explicitly, for a B-spline basis $\{b_{k,r}(f,\bm{\xi})\}_{k=1}^{K}$ of degree $r$ defined on an arbitrary knot sequence $\boldsymbol{\xi}$, the roughness penalty follows the standard Tikhonov regularisation framework and is given by the Gram matrix of the derivatives of the basis functions,
$P_{kl}=\int_{f_{min}}^{f_{max}}b_{k,r}^{\prime}(f;\bm{\xi})\,b_{l,r}^{\prime}(f;\bm{\xi})\,\mathrm{d}f,\qquad k,l=1,\ldots,K.$ (12)
This corresponds to a first-order roughness penalty, i.e., penalisation of the squared $L^{2}$ norm of the first derivative of the spline $b_{k,r}^{\prime}(f,\bm{\xi})$. In our implementation, this matrix is evaluated directly by applying the linear differential operator to the B-spline basis, as provided by skfda [scikit-fda-repo, scikit-fda-paper]. The roughness penalty $\boldsymbol{\lambda}^{\top}\mathbf{P}\boldsymbol{\lambda}$ therefore discourages large local fluctuations in adjacent spline coefficients, enforcuing smoothness in the estimated spectrum.
The rate parameters $\beta_{\phi}$ and $\beta_{\delta}$ are given small prior values (e.g. $10^{-4}$), and the shape parameters $\alpha_{\phi}$ and $\alpha_{\delta}$ are set near 1 for robustness. These weakly informative Gamma priors are standard in hierarchical smoothing models, providing scale-invariant regularisation that encourages stable yet flexible behaviour across different datasets [JULLION2007, Bremhorst2013, maturana-russel_bayesian_2021]. The hierarchical structure decouples the penalty strength from a fixed hyperparameter choice: the latent precision parameter $\phi$ governs the smoothness of the spline coefficients $\boldsymbol{\lambda}$, while the hyperparameter $\delta$ acts as a scaling prior that adaptively regularizes $\phi$. Larger $\phi$ values induce smoother spectra, whereas smaller $\phi$ values permit greater local flexibility. This formulation prevents over- or under-smoothing that could arise from arbitrarily chosen fixed penalty values.
@@
We use a blocked Gibbs sampler to sample from the joint posterior $p(\boldsymbol{\theta}|\textbf{Y})$ by cyclically sampling from each of the full conditional posterior distributions.The parameters $\phi$ and $\delta$ can be sampled directly from their full conditional posterior distributions
@@
Within the Gibbs-sampling framework, $\boldsymbol{\lambda}$ is sampled either using a Metropolis–Hastings method for component-wise updates [maturana-russel_bayesian_2021] or an adaptive Metropolis–Hastings (AMH) algorithm for joint updates [roberts_examples_2009].
For computational efficiency, the knots are placed at the beginning of the algorithm and remain fixed throughout. To ensure sufficient flexibility, we can use a large $N_{\xi}$, as illustrated in the Application section, so that the model can capture the underlying variations of the PSD while the penalty prior guards against overfitting.